troydecpeticon:

homework help please. i copy pasted this from the website: Let f(x) = (x^3 + 2x^2 - 3x + 1) / (x^2 - 1). a) Find the domain of the function f(x). b) Find the vertical asymptotes, if any. c) Find the x-intercepts, if any. d) For which values of x does f(x) have horizontal tangents? e) Calculate the derivative of f(x). f) Find the critical points of f(x). g) Determine the concavity of f(x) and identify any inflection points. h) Sketch the graph of f(x), showing key features such as intercepts, asymptotes, and critical points.

Sandy99:

Damn math theses days ain't the same when I was in school

troydecpeticon:

oh dang

Sowny17:

Is the answer going to be an number or A to H?

Aratox:

This explanation will be long, please bear with me. Your key points are summarized at the bottom. a) To find the domain of the function f(x), we need to determine the values of x for which the function is defined. The function f(x) is defined for all real numbers except those that would make the denominator, x^2 - 1, equal to zero. So, we need to solve the equation x^2 - 1 = 0. x^2 - 1 = 0 (x - 1)(x + 1) = 0 x = 1, x = -1 Therefore, the domain of the function f(x) is all real numbers except x = 1 and x = -1. Domain: (-∞, -1) U (-1, 1) U (1, ∞) b) To find the vertical asymptotes, we need to determine the values of x for which the function approaches positive or negative infinity. As x approaches positive infinity, the function f(x) approaches positive infinity. Therefore, there is no vertical asymptote as x approaches positive infinity. As x approaches negative infinity, the function f(x) approaches negative infinity. Therefore, there is a vertical asymptote at x = -1. Vertical asymptote: x = -1 c) To find the x-intercepts, we need to solve the equation f(x) = 0. (x^3 + 2x^2 - 3x + 1) / (x^2 - 1) = 0 Since the numerator is a cubic polynomial, it does not equal zero for any real values of x. Therefore, the function f(x) does not have any x-intercepts. x-intercepts: None d) To find the values of x for which f(x) has horizontal tangents, we need to find the values where the derivative of f(x) is zero. To calculate the derivative of f(x), use the Quotient Rule: f'(x) = [(x^2 - 1)(3x^2 + 4x - 3) - (x^3 + 2x^2 - 3x + 1)(2x)] / (x^2 - 1)^2 Simplifying the expression: f'(x) = (3x^4 + 6x^3 - 8x^2 - 3x + 3) / (x^2 - 1)^2 Set f'(x) = 0 and solve for x: 3x^4 + 6x^3 - 8x^2 - 3x + 3 = 0 This equation is difficult to solve analytically, but you can use numerical methods, such as graphing or approximation techniques, to find the values of x where f(x) has horizontal tangents. e) The derivative of f(x) is f'(x) = (3x^4 + 6x^3 - 8x^2 - 3x + 3) / (x^2 - 1)^2. f) To find the critical points of f(x), we need to find the values of x for which the derivative f'(x) is either zero or undefined. Since we already found that f'(x) = 0 at some values, those values are the critical points: Critical point: x = (the solutions from step d) Note: If f'(x) were undefined at any x-values, those x-values would also be critical points, but in this case, the derivative is defined for all x-values. g) To determine the concavity of f(x) and identify any inflection points, we need to find the values of x where the second derivative of f(x) changes sign. To calculate the second derivative of f(x), differentiate the first derivative f'(x): f''(x) = [12x^3 + 18x^2 - 16x - 3(x^2 - 1)(3x^2 + 4x - 3)] / (x^2 - 1)^3 Simplifying the expression: f''(x) = (12x^3 + 18x^2 - 16x - 3x^4 - 3x^3 - 6x^2 + 9x + 3) / (x^2 - 1)^3 Set f''(x) = 0 and solve for x to find the potential inflection points. This equation is again difficult to solve analytically, so you can use numerical methods to find the approximate values. h) To sketch the graph of f(x), you can use all the information gathered so far. Key features to include in the graph: - Domain: (-∞, -1) U (-1, 1) U (1, ∞) - Vertical asymptote: x = -1 - X-intercepts: None - Critical points: x = (the solutions from step d) - Concavity and inflection points: Plot the inflection points found in step g. It is also helpful to plot a few additional points corresponding to x-values within the domain to better understand the behavior of the function.